# Understanding and Characterizing Obfuscated Funds Transfers in Ethereum Smart Contracts

### v1

@@ Abstract.
Scam contracts on Ethereum have rapidly evolved alongside the rise of DeFi and NFT ecosystems, utilizing increasingly complex code obfuscation techniques to avoid early detection. This paper systematically investigates how obfuscation amplifies the financial risks of fraudulent contracts and undermines existing auditing tools. We propose a transfer-centric obfuscation taxonomy, distilling seven key features, and design ObfProbe, a framework that performs bytecode-level smart contract analysis to uncover obfuscation techniques and quantify the level of obfuscation complexity via Z-score ranking. In a large-scale study of 1.03 million Ethereum contracts, we isolate over 3,000 highly obfuscated contracts and discover two scam types, three high-risk contract types, and MEV bots, that are deeply coupled with various obfuscation maneuvers such as assembly usage, dead code, and deep function splitting. We further reveal that obfuscation substantially increases the scale of financial damage and evasion time. Finally, we evaluate SourceP, a state-of-the-art Ponzi detection tool, on both obfuscated and non-obfuscated samples, observing its accuracy to drastically fall from  80% (non-obfuscated) to  12% (obfuscated) in real-world scenarios. These findings underscore an urgent need for enhanced “anti-obfuscation” analysis techniques and broader community collaboration to mitigate scam contract proliferation in the expanding DeFi ecosystem.
1. Introduction
@@
Smart contracts, which are self-executing programs deployed on blockchains, have been widely adopted recently. It has enabled various significant emerging applications, such as decentralized finance (Compound, 2025; Burgerswap, 2025; Uniswap, 2025) and digital art trading (Ruan, 2022). Along with it, security issues are also on the rise. Recent reports (Wu et al., 2023; He et al., 2023) show that malicious smart contracts, including scams and MEV bots, have become increasingly prevalent, resulting in significant financial losses. To address these significant threats, researchers have proposed a variety of techniques to detect (Tann et al., 2018; Sayeed et al., 2020; Sathvik and Mazumdar, 2024) and analyze (Gupta et al., 2022b; Shah et al., 2023) these emerging security issues. These techniques rely mainly on static program analysis (Ivanov et al., 2023) and rule-based matching (Agarwal et al., 2022), which have been proven to be highly effective and efficient in detecting early malicious smart contracts, which are typically straightforward and easily identifiable, such as transfers to externally owned private account addresses.
@@
As this arms race continues, malicious smart contracts are gradually replaced by more covert, complex, and obfuscated contract logic (Ano, 2020). Recent studies (Xia et al., 2025; Zhang et al., 2023) confirm that obfuscation has become the primary means by which attackers conceal genuine malicious transfer or backdoor control logic, splitting functions, adding redundant instructions, using assembly code, and other obfuscation techniques. When attackers employ these obfuscation techniques in their malicious smart contracts, the static analysis and matching rules adopted by traditional detection tools are often disrupted, leading to high detection inaccuracies (Sendner et al., 2023; Zhou et al., 2024) and further worsening financial losses (Zhou et al., 2023; Wang et al., 2024; Victor, 2022; Gupta et al., 2022a).
@@
Although obfuscation techniques are witnessed to be increasingly employed in smart contracts to evade security analysis and have been linked to substantial financial losses, no comprehensive study has yet been conducted to systematically assess their impact in real-world scenarios — an understanding that is critical for the development of effective defense mechanisms and the mitigation of associated security risks. In this paper, we report the first systematic study on the obfuscation of funds transfer operations, which are the most essential and security-critical activities of a smart contract (Pan et al., 2023).
Specifically, we aim to thoroughly understand the status quo of obfuscated funds transfer operations in Ethereum smart contracts by answering the following four essential research questions:
@@
To answer the aforementioned research questions, we systematically develop a taxonomy to define and characterize different obfuscation techniques on funds transfer operations and propose ObfProbe, an EVM bytecode analysis tool that can accurately uncover how different obfuscation techniques manifest in real-world smart contracts. With the help of this analysis framework, we conduct a series of comprehensive studies on a huge real-world dataset of 1,042,928 smart contracts.
Here we highlight some interesting discoveries:
@@
    Our analysis framework ObfProbe identified 3,128 highly obfuscated contracts in the wild. Within them, 153 contracts harbor substantial security risks, placing user funds totaling close to \$100 million at risk.
@@
    Compared to non-obfuscated samples, obfuscated scam contracts demonstrate a significantly larger financial impact, with their highest recorded inbound funds being roughly 2.4 times greater and clear periods of intensified victimization occurring between 2019 and 2023.
@@
    Deep obfuscation can seriously undermine existing detection mechanisms (e.g., the accuracy of a state-of-the-art Ponzi detector drops from 79% to 12% when applied to obfuscated code).
@@
  • We developed ObfProbe, the first EVM bytecode obfuscation analyzer leveraging seven bytecode-level features and a Z-score representation model to detect obfuscated transfer logic automatically.
  • We performed the first large-scale measurement study on more than 1.04 million Ethereum smart contracts, revealing that 0.3% exhibit complicated obfuscation and categorizing three major threat contexts: MEV bots, traditional scam tokens, and highly centralized permission wrappers.
  • We quantified the impact of obfuscation on financial damage and detection: obfuscated contracts can extract up to 201.74 ETH per scam and trigger significant victim outbreaks, while state-of-the-art detectors suffer a decrease in accuracy from 79% to under 12% under deep obfuscation.
@@
  • We open-sourced the implementation of our prototype and all study data/artifacts to facilitate future research [^1] https://github.com/dcszhang/Obfuscation_Tool.
2. Background
2.1. Blockchain and Smart Contracts
@@
Blockchain, a technology with a decentralized distributed ledger at its core, ensures data security and immutability through cryptographic methods, making it a foundational infrastructure for various data transactions (Ali and Sofi, 2021). Its core features—decentralization, transparency, immutability, and security—position it as a transformative technological innovation across a wide range of industries (Tariq et al., 2019).
@@
Smart contracts are programs built on top of blockchains that autonomously execute contractual terms when predefined conditions are satisfied, eliminating the need for intermediaries (Rashid and Siddique, 2019). Smart contracts offer several key advantages, including reduced transaction costs, enhanced efficiency, and a lower risk of human error. By permanently recording execution results and data on the blockchain, they eliminate the possibility of unilateral alteration, thereby ensuring fairness and transparency in contract execution (Sakib, 2024). Smart contracts have been widely applied in various fields, including financial transactions, identity verification, supply chain management, and insurance (Sharma and Dutta, 2018). In the financial sector, smart contracts facilitate automatic settlement and payment on the blockchain, significantly enhancing transaction efficiency and transparency while reducing the need for intermediaries and human involvement. As blockchain technology evolves, smart contracts will play an increasingly important role across multiple industries (Tariq et al., 2019).
2.2. Code Obfuscations
Code obfuscation is the process of transforming a program into an equivalent version that preserves its semantics but significantly impedes human understanding and automated analysis. Originating in the late 1990s as a software‐protection technique, obfuscation now finds applications both in legitimate software‐protection (e.g. DRM, piracy resistance) and in malware‐evasion (hindering static and dynamic analysis). While no obfuscation can guarantee unbreakable protection, carefully designed transformations raise the attacker’s cost—forcing reverse engineers to invest more time, specialized tools, or human effort to recover the original logic.
Research over the past years continues to refine a taxonomy of obfuscation techniques, typically grouped into control-flow obfuscation, data obfuscation, layout (lexical) obfuscation, and instruction substitution. Control-flow obfuscation aims to alter the program’s control-flow graph (CFG) without changing semantics, employing opaque predicates, bogus branches, control-flow flattening (e.g. dispatcher loops) or virtualization-based transformations that compile sensitive code into custom bytecode (e.g. VMProtect) (Li et al., 2022; Schloegel et al., 2022). Data obfuscation focuses on concealing computation logic by encoding variables, splitting values, or applying mixed Boolean-arithmetic (MBA) expressions that rewrite simple operations into complex algebraic equivalents—a method effective against pattern-based decompilers and symbolic engines but sometimes simplified by compiler optimizations (Little and Xu, 2023). Layout obfuscation modifies non-functional code aspects such as renaming identifiers, stripping or reordering symbols and functions, and inserting NOPs or dummy code; albeit low-cost, such lexical changes provide limited protection if used alone (Collberg et al., 1997). Instruction substitution replaces code sequences with semantically equivalent but syntactically distinct alternatives— for instance, unrolling multiplications into shift-add sequences or recasting boolean logic into arithmetic forms—to diversify byte patterns and defeat signature-based analysis and metamorphic malware detection (Schloegel et al., 2022).
3. Taxonomy of Obfuscated Funds Transfers
@@
In this paper, we primarily focus on funds transfer operations, which are the most essential and security-critical activities of a smart contract (Pan et al., 2023). To answer RQ1 and further conduct our study, we develop a taxonomy to define and characterize different obfuscation techniques on funds transfer operations, based on how each component of funds transfer operations can be hidden, derived from our deep domain knowledge of Ethereum smart contracts.
@@
Our taxonomy development starts by dissecting every element of the funds transfer operation. The standard way to realize such an operation is through a transfer API CALL in Solidity (Solidity, 2025), which is implemented as a CALL EVM opcode in a given smart contract bytecode. Specifically, we define a CALL EVM opcode that implements a funds transfer operation as follows:
@@
(1) $T=(addr,value,context,log),where:\vspace{-0.05in}$
@@
$\bullet$ context is the execution context of the funds transfer operation, which includes the storage state of the contract, the remaining instructions and control/data flow inside the function where the transfer is located. It together determines whether and how the CALL instruction can be executed.
@@
Based on our definition of a funds transfer operation, we investigate each element and derive seven obfuscation strategies. By exhaustively mapping each strategy to one or more elements of a funds transfer operation, we ensure our taxonomy is comprehensive and covers all fundamental ways to hide transfer operations in Ethereum smart contracts.
Please note that the source code listings below are for illustration purposes, while our analysis is done at the bytecode level.
3.1. Obfuscation of addr
The addr element, which is essentially a string that represents a 20-byte Ethereum address, indicates the recipient address of a funds transfer operation. We consider four obfuscation methods, stemming from traditional string obfuscation techniques (Quiring et al., 2019).
@@ 3.1. Obfuscation of addr
Listing 1: T1. Multi-step Address Generation
@@
Listing 2: T2. Complex String Operations
T3. External Contract Calls. Instead of local computation, the contract with obfuscation may choose to query a “router” or “delegate” contract to fetch addr, hiding the true recipient behind an external CALL.
@@
2function getAddr() external returns (address);
@@
Listing 3: T3. External Contract Calls
T4. Control‐flow complexity. Dynamic conditions (e.g., block.timestamp) select among multiple branches—some dead—each leading to different addresses. This conditional complexity obscures which branch yields the actual addr.
@@
14}
15rec.transfer{value: value}("");
Listing 4: T4. Control-flow Complexity
3.2. Obfuscation of value
Since value is the amount transferred in wei, i.e., a 256-bit unsigned integer. Therefore, two obfuscation strategies from addr can be applied.
@@ 3.3. Obfuscation of context
4}
5// no-op arithmetic
6uint x = (1 + 2) - 3;
7// core transfer hidden among noise
8address rec = 0xAbCd...;
9rec.transfer{value,...}("");
Listing 5: T5. Camouflage Instructions
T5. Camouflage Instructions. By injecting large numbers of meaningless loops, arithmetic operations, and No-ops into the transfer-related function body, the core CALL is camouflaged with many irrelevant instructions. Although the transfer is still executed correctly, the altered control and data flow within the function make it hard for static analysis tools to isolate the actual CALL context.
@@
Listing 6: T6. Replicated Transfer Logic
3.4. Obfuscation of log
@@
T7. Irrelevant Log Events. One can emit misleading or unrelated events (e.g., logging a transfer to a “legitimate” address while sending funds elsewhere), diverting auditors’ and tools’ attention, concealing the real log data that corresponds to the transfer.
@@
Listing 7: T7. Irrelevant Log Events
4. System Design and Implementation
Building on the taxonomy, we design and implement an EVM bytecode analysis tool named ObfProbe to uncover how different obfuscation techniques manifest in real-world smart contracts so as to answer our RQ2. In this section, we elaborate on the design and implementation of our system.
4.1. System Overview
Figure 1. Overview of ObfProbe.
Figure 1 shows an overview of ObfProbe. At a high level, it first converts a given smart contract bytecode to the static single assignment intermediate representation (SSA IR) using an existing tool named Rattle (Crytic, 2025). After that, it scans the IR to detect all funds transfer operations (stage 1). For each transfer, our system extracts seven pre-defined obfuscation features from the SSA IR (stage 2). Finally, it applies a Z-score representation model (Curtis et al., 2016), which is a numerical value that represents a data point’s distance from the mean in terms of standard deviations, to convert the extracted features to an obfuscation score (stage 3), indicating the degree of obfuscation applied to the transfer operation.
Table 1. Summary of Transfer-related Obfuscation Features

### v2

@@ Abstract
Scam contracts on Ethereum have rapidly evolved alongside the rise of DeFi and NFT ecosystems, utilizing increasingly complex code obfuscation techniques to avoid early detection. This paper systematically examines how obfuscation exacerbates the financial risks associated with fraudulent contracts and undermines existing auditing tools. We propose a transfer-centric obfuscation taxonomy, distilling seven strategies and seven measurable signals, and design ObfProbe, a framework that performs bytecode-level smart contract analysis to uncover obfuscation techniques and quantify the level of obfuscation complexity via Z-score ranking. In a large-scale study of 1.04 million Ethereum contracts, we isolate over 3,000 highly-obfuscated contracts and, in manual case studies, identify four patterns: MEV bots, Ponzi schemes, fake decentralization, and extreme centralization that are deeply coupled with various obfuscation maneuvers, including assembly usage, dead code, and deep function splitting. We further reveal that obfuscation substantially increases the scale of financial damage and the time required for evasion. Finally, we evaluate SourceP, a state-of-the-art Ponzi detection tool, on both obfuscated and non-obfuscated samples, observing recall dropping from 0.80 to 0.12 and F1 from 0.89 to 0.21 under deep obfuscation in real-world scenarios. These findings underscore the urgent need for enhanced “anti-obfuscation” analysis techniques and broader community collaboration to mitigate the proliferation of scam contracts in the expanding DeFi ecosystem.
Index Terms:
smart contracts, obfuscation, Ethereum, Ponzi, MEV, static analysis
I Introduction
@@
Smart contracts, which are self-executing programs deployed on blockchains, have been widely adopted recently. It has enabled various significant emerging applications, such as decentralized finance [12, 9, 64] and digital art trading [47]. Along with it, security issues are also on the rise. Recent reports [68, 26] show that malicious smart contracts, including scams and MEV bots, have become increasingly prevalent, resulting in significant financial losses. To address these significant threats, researchers have proposed various techniques to detect [60, 50, 49] and analyze [23, 55] these emerging security issues. Many of the techniques rely on static program analysis [27] and rule-based matching [1], which have been proven highly effective and efficient in detecting early malicious smart contracts, typically straightforward and easily identifiable, such as transfers to externally owned private account addresses.
@@
As this arms race continues, malicious smart contracts are gradually replaced by more covert, complex, and obfuscated contract logic [69, 71, 54]. Recent studies [69, 71] confirm that obfuscation has become the primary means by which attackers conceal malicious transfer or backdoor control logic, which includes the use of assembly code, splitting functions, and redundant instructions. When attackers employ such obfuscation techniques in their smart contracts, the static analysis and matching rules adopted by traditional detection tools are often disrupted, leading to high detection inaccuracies [54, 72] and further exacerbating financial losses [73, 67, 65, 24]. Moreover, obfuscation techniques are not only found in malicious contracts but also in benign contracts for various reasons (e.g., to protect proprietary business logic and deter copycat attacks), which hinders security analysis and prevents users from better understanding the behaviors of the contracts.
@@
Although obfuscation techniques are increasingly employed in smart contracts and have been linked to substantial financial losses, no comprehensive study has yet systematically assessed their impact in real-world scenarios, an understanding that is critical for developing effective defense mechanisms and mitigating associated security risks. In this paper, we aim to conduct the first systematic study on the obfuscation of funds transfer operations, which are the most essential and security-critical activities of a smart contract [42]. In particular, we aim to thoroughly understand the status quo of obfuscated funds transfer operations in Ethereum smart contracts by answering the following four essential research questions:
@@
To answer these research questions, we develop a taxonomy to define and characterize obfuscation techniques in fund transfer operations. Specifically, we propose seven robust features by systematically dissecting and examining the formal definition of funds transfer operations within the Ethereum virtual machine to ensure the comprehensiveness and the representativeness of the feature list. We quantify the obfuscation complexity of fund transfer operations in smart contracts using a Z-score-based representation model. We then propose ObfProbe, an EVM bytecode analysis framework that accurately uncovers various obfuscation techniques used in real-world smart contracts. Using this analysis framework, we conduct a series of studies on real-world smart contracts to examine the use of obfuscation techniques for both malicious and benign purposes in funds transfer operations. Below, we highlight some interesting discoveries:
@@
    By analyzing the top 3000 highly obfuscated contracts detected by ObfProbe, we found 463 contracts that exhibit substantial security risks, placing funds totaling $\approx$\$100 million at risk.
@@
    Compared to non-obfuscated scam contracts, obfuscated scam contracts demonstrate a significantly larger financial impact, with their highest recorded inbound funds being $\approx$2.4X higher and clear periods of intensified victimization occurring between 2018 and 2025.
@@
    The evaluation of a state-of-the-art Ponzi detector on obfuscated Ponzi contracts shows a significant effectiveness downgrade (e.g., F1 score from 0.89 to 0.21), demonstrating that obfuscations can significantly undermine the performance of existing detection tools.
@@
  • We developed ObfProbe, the first EVM bytecode obfuscation analyzer that leverages seven bytecode-level features and a Z-score representation model to automatically detect obfuscated transfer logic from smart contracts.
  • We conducted a large-scale measurement of more than 1.04 million Ethereum smart contracts, revealing common obfuscation patterns in the top 3000 contracts, including four types: MEV bots, Ponzi schemes, fake decentralization, and extreme centralization.
  • We quantitatively studied the impact of obfuscation on financial damage and detection.
@@
  • We open-sourced ObfProbe and the collected data/artifacts to facilitate future research [^1] https://github.com/nonname-byte/Obfuscation_Tool.
II Background
II-A Blockchain and Smart Contracts
@@
Blockchain, a technology with a decentralized distributed ledger at its core, ensures data security and immutability through cryptographic methods, making it a foundational infrastructure for various data transactions [2]. Its core features—decentralization, transparency, immutability, and security—position it as a transformative technology across a wide range of industries [61].
@@
Smart contracts are programs built on top of blockchains that autonomously execute contractual terms when predefined conditions are satisfied, eliminating the need for intermediaries [46]. Smart contracts offer several key advantages, including reduced transaction costs, enhanced efficiency, and a lower risk of human error. By permanently recording execution results and data on the blockchain, they eliminate the possibility of unilateral alteration, thereby ensuring fairness and transparency in contract execution [48]. Smart contracts have been widely applied in various fields, including financial transactions, identity verification, supply chain management, and insurance [48, 61]. In the financial sector, smart contracts enable the automatic settlement and payment on the blockchain, significantly enhancing transaction efficiency and transparency while reducing the need for intermediaries and human intervention. As blockchain technology evolves, smart contracts will play an increasingly important role across multiple industries [61].
II-B Code Obfuscations
Code obfuscation transforms a program into a semantics-preserving but harder-to-analyze form, used both for software protection and for malware evasion to raise reverse-engineering cost. A standard taxonomy groups techniques into control-flow, data, layout (lexical), and instruction-substitution transformations [11]. Control-flow obfuscation perturbs the CFG (e.g., opaque predicates, bogus branches, flattening) and even virtualization to hinder static/dynamic analyses [33, 52]. Data obfuscation hides computations via variable encoding and mixed Boolean–arithmetic (MBA) rewriting; layout changes rename/reorder identifiers or inject dead/NOP code; instruction substitution replaces code with equivalent forms (e.g., shifts/adds for multiply), which diversifies byte patterns and weakens signature-based detection [36, 11, 52].
III Taxonomy of Obfuscated Funds Transfers
@@
Our study focuses on funds transfer operations, which are the most essential and security-critical activities of a smart contract [42]. To answer our research questions, we develop a taxonomy to define and characterize different obfuscation techniques on funds transfer operations, based on how each component of funds transfer operations can be hidden, derived from the formal definition of funds transfer operations within the Ethereum virtual machine to ensure the comprehensiveness and the representativeness of our taxonomy.
@@
We start by dissecting every element of the funds transfer operation. The standard way to realize such an operation is through a transfer API CALL in Solidity [58], which is implemented as a CALL EVM opcode in a given smart contract bytecode. Specifically, we define a CALL EVM opcode that implements a funds transfer operation as follows:
@@
$T=(addr,value,context,log),where:\vskip-3.61371pt$ (1)
@@
$\bullet$ context is the execution context of the funds transfer operation, which includes the storage state of the contract, the remaining instructions, and control/data flow inside the function where the transfer is located. This information together determines whether and how the CALL instruction can be executed.
@@
Based on our definition of a funds transfer operation, we investigate each element and derive seven obfuscation features. By exhaustively mapping each feature to one or more elements of a funds transfer operation, we ensure our taxonomy is comprehensive and covers all fundamental methods of hiding transfer operations in Ethereum smart contracts. Please note that our analysis is conducted at the bytecode level, and the source code listings below are provided for illustrative purposes.
III-A Obfuscation of addr
The addr element is essentially a string that represents a 20-byte Ethereum address, indicating the recipient address of a funds transfer operation. We consider four obfuscation methods that stem from traditional string obfuscation techniques [45].
@@ III-A Obfuscation of addr
Listing 1: T1 Multi-step Address Generation
@@
Listing 2: T2 Complex String Operations
T3. External Contract Calls. Instead of local computation, the contract with this obfuscation technique may choose to query a “router” or “delegate” contract to fetch addr, hiding the true recipient behind an external CALL.
@@
2 function getAddr() external returns (address);
@@
Listing 3: T3 External Contract Calls
T4. Control-flow complexity. The branch selection is dependent on run-time conditions (e.g., block.timestamp). Some branches are dummy and never execute, and different branches point to different addrs. This hides the real recipient because the true branch is known only at execution time, which hinders static rule matching.
@@
13rec.transfer{value: value}("");
Listing 4: T4 Control-flow Complexity
III-B Obfuscation of value
Since value is the amount transferred in wei, i.e., a 256-bit unsigned integer. Therefore, two obfuscation strategies from addr can also be applied.
@@ III-C Obfuscation of context
T5. Camouflage Instructions. By injecting large numbers of meaningless loops, arithmetic operations, and NOPs into the transfer-related function body, the core CALL is camouflaged with many irrelevant instructions. Although the transfer is still executed correctly, the altered control and data flow within the function make it hard for static analyzers to isolate the actual context.
@@
4// no-op arithmetic
5uint x = (1 + 2) - 3;
6// core transfer hidden among noise
7address rec = 0xAbCd...;
8rec.transfer{value,...}("");
Listing 5: T5 Camouflage Instructions
@@
Listing 6: T6 Replicated Transfer Logic
III-D Obfuscation of log
@@
T7. Irrelevant Log Events. One can emit misleading or unrelated events (e.g., logging a transfer to a “legitimate” address while sending funds elsewhere), diverting auditors’ and tools’ attention, and concealing the real log data that corresponds to the transfer.
@@
Listing 7: T7 Irrelevant Log Events
Answer to RQ1. We answer RQ1 by formalizing the EVM transfer path, deriving seven obfuscation patterns (T1–T7) and mapping each to measurable bytecode features (F1–F7). This taxonomy operationalizes transfer obfuscation and provides the quantitative basis used by our Z-score model in Section IV.
IV System Design and Implementation
Building on the taxonomy, we design and implement an EVM bytecode analysis tool named ObfProbe to uncover how different obfuscation techniques manifest in real-world smart contracts, thereby enabling our RQ2 measurement.
IV-A System Overview
Figure 1: Overview of the analysis pipeline of ObfProbe.
Figure 1 shows an overview of ObfProbe. At a high level, it converts a given smart contract bytecode to the static single assignment intermediate representation (SSA IR) using an existing tool named Rattle [15]. After that, it scans the IR to detect all fund transfer operations. For each transfer, our system extracts seven pre-defined obfuscation features. Finally, it applies a Z-score representation model [16], which is a numerical value that represents a data point’s distance from the mean in terms of standard deviations, to convert the extracted features into an obfuscation score that indicates the degree of complexity in the obfuscation applied to the transfer operation.
TABLE I: Summary of transfer-related obfuscation features.